OpenStudy (anonymous):

Jeanette can make 22 pompoms in 6 hours. If she works 18 hours, how many pompoms can she make? a. 79.2 pompoms c. 60 pompoms b. 66 pompoms d. 82.67 pompoms

OpenStudy (anonymous):

22/6 = x/18 (22 pompoms in 6 hrs = x pompoms in 18 hrs) cross multiply (6)(x) = (22)(18) 6x = 396 x = 396/6 x = 66 any questions ?

OpenStudy (anonymous):

no thanks for showing the work though(: And thanks for helping me!

OpenStudy (anonymous):

no problem :)